Soybean meal is 18% protein, cornmeal is 9% protein.How many pounds of each should be mixed in order to get 360lb of mixture that is 10% protein?

a. how many pounds of cornmeal should be in the mixture?

b. how many pounds of soybean meal should be in the mixture?

Answer by anantha(86)

let no of pounds of soybean meal be x pounds
and no of pounds of cornmeal be y pounds
x pounds of 18% protein and y pounds of 9% protein should be mixed we get 360lb of 10% protein
x+y=360lb -------(1)equation
y=360-x(i.e; for isolating y subtract x from both sides)
x * (18%)+ y*(9%) =360*(10%)
x*(18/100)+y*(9/100)=360*(10/100)
18x/100+9y/100=3600/100
multiplying every term by 100
(18x/100)*100+(9y/100)*100=(3600/100)*100
after removing 100 on both sides we have
18x+9y=3600------------(2) equation
substitute y value in 2 equation
18x+9(360-x)=3600
18x+9*360-9*x=3600
18x+3240-9x=3600
9x+3240=3600(subtracting 9x from 18x we get 9x)
isolate x on one side subtract 3240 on both sides
9x+3240-3240=3600-3240
9x=360
dividing both sides by 9 we get
9x/9 =360/9
x=40
substitute x value in y=360-x
y=360-x
y=360-40=320
no of pounds of soybean meal=40 pounds
no of pounds of cornmeal=320 pounds